Mark is 3 times as old as Julie. The difference in their ages is 10 years.
What is Mark’s age?

Answer by god2012(113)

Let 'M' and 'J' are the ages of Mark and Julie respectively.
Given :
M = 3J
M - J = 10 or J = M - 10
M = 3 (M+10) = 3M - 30
2M = 30 or M = 15 years